OpenStudy (anonymous):

help 4 medal Questions_ How do quadratic functions compare to linear and exponential functions? How are solutions to a system of equations graphed?

OpenStudy (anonymous):

help

OpenStudy (kirbykirby):

Quadratic function: \(f(x) = ax^2 + bx + c\)... you have an exponent 2 as the highest degree Linear function: \(f(x) = ax + b\), the highest degree is 1 Exponential function \(f(x) = a(b)^x+c\), here "x" is in your exponent

OpenStudy (anonymous):

so thats the answer?

OpenStudy (kirbykirby):

Well part of it.. I am guessing they want you to write sentences to describe them more fully. YOu can also describe how the general shapes of the graphs differ..